OpenStudy (kendricklamar2014):

Solve the following compound equality. 5m < -10 or 3m > 9

OpenStudy (anonymous):

Do you have options?

OpenStudy (kendricklamar2014):

no

OpenStudy (anonymous):

Refer to this, hopefully it can help you. http://www.coolmath.com/algebra/07-solving-inequalities/06-compound-inequalities-01.htm

OpenStudy (kendricklamar2014):

@dumbcow

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5m < -10~~ or~~ 3m>9\\~\\ -\infty\le 5m \le -10~~ or~~ 9\le 3m\le +\infty\\~\\ -\infty\le m \le -2 ~~or ~~3\le m\le +\infty\\~\\ (-\infty,-2)\cup(3,+\infty)}\end{align}\)

OpenStudy (mathmath333):



OpenStudy (mathmath333):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

i dont get this??

OpenStudy (mathmath333):

which part?

OpenStudy (kendricklamar2014):

this part.....